OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 3) and (2, 7)? x - y = -2 x - y = -1 x - y = - 6 x - y = - 5

hartnn (hartnn):

first, can you find the slope between these 2 points (-2, 3) and (2, 7)?

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (anonymous):

I got 1

hartnn (hartnn):

thats correct! :)

hartnn (hartnn):

now, the equation when you have a slope and point is y - y1 = m (x -x1) where m= 1 x1 = -2, y1 = 3 just plug in values!

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

wait, I am confused

hartnn (hartnn):

confused where ? x1 = -2, y1 = 3 ?

OpenStudy (anonymous):

yeah

hartnn (hartnn):

you can use any one of the given 2 points... (x1,y1) = (-2,3) ---->x1 = -2, y1 = 3 or (x1,y1) = (2,7) ---->x1 = 2, y1 = 7

hartnn (hartnn):

just plug any one set in the standard form \(\large y-y_1=m(x-x_1)\) and simplify

OpenStudy (anonymous):

oh okay

hartnn (hartnn):

so, can u get the equation now ?

OpenStudy (anonymous):

yea

hartnn (hartnn):

if you want to verify your equation, tell me what u got...

OpenStudy (anonymous):

I got the last one

hartnn (hartnn):

bravo! you are correct :)

OpenStudy (anonymous):

awesome! Thanks :)

hartnn (hartnn):

welcome ^_^